OpenStudy (anonymous):

solve and check 5(y+6) = 5(y-3)

OpenStudy (anonymous):

5(y+6=5(y-3) 5y+30=5y-3 5y+30+3=5y-3+3 5y+33=5y 5y+33-5y=5y-5y y=33

OpenStudy (anonymous):

5(y+6)=5(y-3) 5y+30=5y-15 +15 +15 5y+45=5y -5y -5y 45=y

OpenStudy (anonymous):

Sorry , she's right I made a mistake somewhere.

OpenStudy (anonymous):

thank you kelsey

OpenStudy (anonymous):

No problem once you get "y" you can plug it back in to see if you got the right answer they should equal each other.

OpenStudy (anonymous):

I was wrong. The y's cancel out, so that means that 45=0 which is FALSE so the answer is, NO SOLUTION.

OpenStudy (anonymous):

@BurrDaBoss 's mistake comes from not distributing correctly.

OpenStudy (anonymous):

5(y+6)=5(y-3) 5y+30=5y-15 add 15 to both sides 5y+45=5y subtract 5y from both sides 45 = 0 impossible no solution @kelsey1 was right